OpenStudy (anonymous):

what is the maximum function of 9+7x-6x^2?

OpenStudy (anonymous):

*max value of...

OpenStudy (anonymous):

take its derivative and you get -12x+7=0 and x=-7/12 plug this into equation

OpenStudy (bahrom7893):

-12x+7=0 -12x=-7 x=7/12, plug that into original

OpenStudy (anonymous):

9-7*7/12-6*49/144

OpenStudy (bahrom7893):

it's positive 7/12

OpenStudy (anonymous):

sorry x=7/12

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

if you don't know what a derivative is, find the vertex

OpenStudy (anonymous):

first coordinate of the vertex is \[-\frac{b}{2a}\] second coordinate is what you get when you replace x by that value, and that will be your maximum

OpenStudy (anonymous):

in fact, this is the method i would use irrespective of whether you know calculus or not

OpenStudy (anonymous):

could you complete the square please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

since \[-\frac{b}{2a}=\frac{7}{12}\] and since the second coordinate is \[\frac{265}{24}\] when you complete the square you will get \[-6(x-\frac{7}{12})^2+\frac{265}{24}\]

OpenStudy (anonymous):

this is by far the easiest way to complete a square when the leading coefficient it not 1

OpenStudy (anonymous):

thanks a lot satellite

OpenStudy (anonymous):

yw